OpenStudy (anonymous):

Which fraction falls between five-eighteenths and five-ninths? one-eighteenth two-ninths four-ninths thirteen-eighteenths

OpenStudy (anonymous):

Put all of them in terms of eighteenths to make it easier to see. You want a fraction between 5/18 and 5/9. 5/9 = 10/18. So you want a fraction equal to between 5 and 10 eighteenths. Your choices are: 1/18 2/9 = 4/18 4/9 = 8/18 13/18 Which one is between 5/18 and 10/18?